Given a population of 5000 score with mean μ=86 and σ=10. How many scores are below 56?




1
Expert's answer
2022-05-22T23:07:08-0400
"P(X<56)=P(Z<\dfrac{56-86}{10})"

"=P(Z<-3)\approx0.00135"

"0.00135\cdot 5000=7"

7 scores are below 56.